An auto travels at rate of 25km/h for 4.0mins, 50km/h at 8.0mins finally at 20km/h at 2.0mins. Find: 1. The total distance covered in km 2. The average speed for the trips in m/s
1
Expert's answer
2018-10-23T11:41:09-0400

The distance covered = [(25 x 4/60) + (50 x 8/60) + (20 x 2/60)] km

= [100/60 + 400/60 + 40/60 ] = 540/60 = 9 km


Average speed = 〖9×10〗^3/((4+8+2) ) m/mins. = 〖9×10〗^3/((4+8+2) x 60) m/sec. = 10.714 m/sec.




Answer: 1. Total distance covered is 9 km .

2. The average speed for the trips is 10.714 m/sec